You own the following stocks in your portfolio. What is the portfolio return?

Stock Invested Amount Return
A 39,388 -4.48
B 43,939 25.40
C 32,229 28.56
D 7,602 15.34
Answer #1
Total Portfolio value = Value of A + Value of B + Value of C + Value of D
=39388+43939+32229+7602
=123158
Weight of A = Value of A/Total Portfolio Value
= 39388/123158
=0.3198
Weight of B = Value of B/Total Portfolio Value
= 43939/123158
=0.3568
Weight of C = Value of C/Total Portfolio Value
= 32229/123158
=0.2617
Weight of D = Value of D/Total Portfolio Value
= 7602/123158
=0.0617
Return of Portfolio = Weight of A*Return of A+Weight of B*Return of B+Weight of C*Return of C+Weight of D*Return of D
Return of Portfolio = -4.48*0.3198+25.4*0.3568+28.56*0.2617+15.34*0.0617
Return of Portfolio = 16.05%